OpenStudy (anonymous):

Write a two-step equation that needs the steps of addition then division in order to be solved, and then solve it. Show your work.

OpenStudy (mertsj):

\[3x-6=15\]

OpenStudy (mertsj):

\[3x-6+6=15+6\]

OpenStudy (mertsj):

\[3x=21\]

OpenStudy (mertsj):

\[\frac{3x}{3}=\frac{21}3\]

OpenStudy (mertsj):

\[x=7\]

OpenStudy (anonymous):

that is correct